a metal weighing 6 mg and containing 4% copper is melted and mixed with 2 mg of a different alloy which contains 80% copper.What percent of the resulting alloy is copper?



Found 3 solutions by Theo, greenestamps, josgarithmetic:
Answer by Theo(13342)

4% * 6 mg = .24 mg of copper in 6 mg of metal 1.
80% * 2 mg = 1.6 mg of copper in 2 mg of metal 2.
total copper is 1.84 mg in total metal of 8 mg.
total copper percent of total metal = 1.84 / 8 * 100 = 23% copper in the combined metal alloy.

Answer by greenestamps(13200)



A fast and easy informal solution, if your mental arithmetic is good....

The fraction of the mixture that contains the higher percentage of copper is 2/(6+2) = 2/8 = 1/4.

That means the percentage of copper in the mixture is 1/4 of the way from 4% to 80%.

From 4% to 80% is a difference of 76%; 1/4 of that is 19%.

The percentage of copper in the mixture is 4%+19% = 23%.

ANSWER: 23%


Answer by josgarithmetic(39617)

Make the accounting!

How much pure copper?
How much of the total metal alloy mixture?
Do what you know to do!